Blue Mountain Corp. currently has 150 million common shares outstanding. In a recent earnings conference call, the CFO estimated that this year’s earnings will be $600 million, and the earnings are expected to grow at a constant rate of 5%. The company’s payout policy is as follows: 30% of the earnings to pay dividends, 20% of the earnings to repurchase stocks, and the rest of the earnings to retain within the firm. If the company’s cost of equity is 14%, what is the stock price per share?
Answer #1

Earning in Year 0 = $600 million

Dividend Paid out in Year 0 = 30% of 600 m = $180 million

Dividend per shared D0 = Total Dividend Paid out / Number of common shares outstanding = 180/150 = $1.2

Growth Rate = g = 5% or 0.05

Dividend in Next period D1 = D0*(1+g) = 1.2*1.05 = $1.26

Cost of Equity r = 14% or 0.14

According to Gordons Growth model,

P0 = D1/(r - g) = 1.26/(0.14 - 0.05) = $14

Hence, current price of share is $14